OpenStudy (anonymous):

Find the general solution of:

OpenStudy (anonymous):

Trigonometry or Differential Equations?

OpenStudy (anonymous):

\[y'''' +2y'''+3y''+2y'+y = 0\]

OpenStudy (anonymous):

So from the characteristic equation I get \[r^4 + 2r^3 + 3r^2 + 2r + 1 = 0\] My problem is I'm not seeing how to factor this?

OpenStudy (anonymous):

Find its characteristics Equation..

OpenStudy (anonymous):

Oh, you have started it..

OpenStudy (amistre64):

if it has rational factors, try descartes rule of sign

OpenStudy (amistre64):

it has no rational roots ....

OpenStudy (amistre64):

what methods are you aware of?

OpenStudy (anonymous):

For factoring?

OpenStudy (amistre64):

i was thinking for finding solutions to diffy qs ... factoring this by hand i would think is pretty much impossible in a reasonable amount of time

OpenStudy (anonymous):

In my class we just started covering nth order linear homogeneous DEQ's

OpenStudy (amistre64):

have you covered an reduction formulas?

OpenStudy (anonymous):

Are you sure, your question is absolutely right?

OpenStudy (amistre64):

my best bet would be to try a power series solution, but that might not be so elegant for this setup

OpenStudy (anonymous):

No we have not covered those yet... hmm does this help at all i Just noticed there is a hint says this equation is \[(r^2+r+1)^2\]

OpenStudy (amistre64):

let\[y=\sum a_nx^n\\ y'=\sum_0 a_n~n~x^{n-1}\\ y''=\sum_1 a_n~n(n-1)~x^{n-2}\\ y'''=\sum_2 a_n~n(n-1)(n-2)~x^{n-3}\\ y''''=\sum_3 a_n~n(n-1)(n-2)(n-3)~x^{n-4}\\ \] substitute, line up your x^k values and adjust your indexes to match; then its just a matter of working out the resulting solutions

OpenStudy (amistre64):

lol, use the hint, by all means :)

OpenStudy (amistre64):

sites going wonky on me ... :(

OpenStudy (amistre64):

r^2 + r + 1 = 0 quad formula r = [-2 +- sqrt(1-4(1)(1))]/2 r = -1 +- sqrt(-3)/2